OpenStudy (anonymous):

Which choice shows equations that can be used to find the area of JKL? (Theres 2 triangles, RST with a base of 18 cm and height of 9 cm, and JKL, with a base of 12 cm and height of x cm.)

OpenStudy (anonymous):

Please explain

OpenStudy (anonymous):

the area of a triangle = 1/2 * base * height

Parth (parthkohli):

Do the triangles have the same areas?

OpenStudy (anonymous):

for jkl base = 12 and height = h

OpenStudy (anonymous):

no, RST has a height of 9 cm and a base of 18 cm, JKL has a height of unknown cm and base of 12 cm.

OpenStudy (anonymous):

* height = x

Parth (parthkohli):

1/2 * Base * Height is the area: 6x is the area

OpenStudy (anonymous):

I think i know the answer, H, first use 9/18 =12/h and then use are= 1/2 (12h)

OpenStudy (anonymous):

sorry my pc is slow

Parth (parthkohli):

Yeah, that's exactly what I stated. 6x(or 6h) is the area. :)

OpenStudy (anonymous):

So thats the answer? tnx, Parth!

Parth (parthkohli):

My pleasure :)

OpenStudy (anonymous):

:D